OpenStudy (anonymous):

der r 4 set of weights (1kg, 2kg, 4kg 8kg 16kg 32kg) the set is good enough to wt any object having an integral wt b/w 1kg and 63kg eg: 19kg=16kg+2kg+1kg if 4kg wt is lost how many wts b/w 1kg and 63kg can no longer be measured

OpenStudy (anonymous):

ans is 32

OpenStudy (anonymous):

Thinking it backwards, you have six weights, and you can choose them to form different weights. i.e., choose one and two, and you get three. choose 32 and 1 and you get 33. You can choose or don't choose the weights to form the weights. So, with six weights, ways to choose is (choose) or (not choose)for each weight thus, is 2*2*2*2*2*2=64 with 4 gone, we are left with 5 weights. thus, similarly, 2*2*2*2*2=32 ANS is 64-32=32

OpenStudy (anonymous):

@Shadowys thanks

OpenStudy (anonymous):

You're welcome!